# On Random Fields Associated with Analytic Wavelet Transform

### v1

@@ 3.3. Covariance of AWT’s magnitudes and phases
Corollary 7 shows that the elements of $\mathbf{A}$ and $\mathbf{B}$ are pairwise independent. It also follows that the random vectors $\mathbf{A}$ and $\mathbf{B}$ are uncorrelated; that is, their $m\times n$ covariance matrix is the zero matrix. However, this does not imply that $\mathbf{A}$ and $\mathbf{B}$ are jointly independent because they are not Gaussian distributed. This finding can be regarded as an initial step toward establishing statistical behavior of the quantifications of phase-amplitude coupling and phase-phase coupling [12, 13]. Choose $m=n$, times $t_{j}=t_{m+j}=j\Delta$ for all $j=1,\ldots,m$, where $\Delta$ is a discretization period, and scales so that $s_{1}=\ldots=s_{m}$ much larger than $s_{m+1}=\ldots=s_{2m}$. Under this configuration, $\mathbf{A}$ and $\mathbf{B}$ can be interpreted as the phase and amplitude, respectively, of bandpass-filtered signals. The uncorrelation between $\mathbf{A}$ and $\mathbf{B}$ when the input is pure noise highlights the need for caution when interpreting indices designed to quantify phase-amplitude and phase-phase coupling [12, 13], particularly when higher order moments are involved. A more thorough investigation of this issue will be pursued in future work.
@@ 4. Conclusion
This paper presented a statistical framework for analyzing the magnitude and phase of the AWT of noisy signals. Modeling the observation as a deterministic signal corrupted by stationary Gaussian noise, we derived joint and marginal probability density functions for the magnitude and phase in the general and noise-only settings. The analysis yielded SNR-dependent concentration bounds, explicit covariance expressions, and independence results, offering mathematical foundations for their distributional and dependence structures. We further connected these results to practical tasks by characterizing ridge detection error probabilities, the regularity of scalogram contours and the relationship between the magnitude and phase. The results lay the groundwork for developing rigorous AWT-based algorithms and wavelet-domain statistical inference.
@@ 5.1. Proof of Lemma 1
(49) $(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}\Gamma(\ell,\ell^{\prime})$
$=$ $(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}\mathbb{E}\left[W_{\Phi}(t_{\ell},s_{\ell})\overline{W_{\Phi}(t_{\ell^{\prime}},s_{\ell^{\prime}})}\right]$
@@
$=$ $\int_{\mathbb{R}}\int_{\mathbb{R}}e^{it_{\ell}\lambda}\overline{\widehat{\psi}(s_{\ell}\lambda)}e^{-it_{\ell^{\prime}}\lambda^{\prime}}\widehat{\psi}(s_{\ell^{\prime}}\lambda^{\prime})\mathbb{E}\left[Z_{F}(d\lambda)\overline{Z_{F}(d\lambda^{\prime})}\right].$
@@
(50) $(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}C_{\ell\ell^{\prime}}$
$=$ $(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}\mathbb{E}\left[W_{\Phi}(t_{\ell},s_{\ell})W_{\Phi}(t_{\ell^{\prime}},s_{\ell^{\prime}})\right]$
@@
$=$ $\int_{\mathbb{R}}\int_{\mathbb{R}}e^{it_{\ell}\lambda}\overline{\widehat{\psi}(s_{\ell}\lambda)}e^{it_{\ell^{\prime}}\lambda^{\prime}}\overline{\widehat{\psi}(s_{\ell^{\prime}}\lambda^{\prime})}\mathbb{E}\left[Z_{F}(d\lambda)Z_{F}(d\lambda^{\prime})\right].$
@@ 5.2. Proof of Proposition 1
$p\left(r;|W_{Y}(t,s)|\right)$
$=$ $\frac{r}{\pi\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\int_{-\pi}^{\pi}\textup{exp}\left(-\frac{\left|re^{i\theta}-W_{f}(t,s)\right|^{2}}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\right)d\theta$
@@
$p\left(r;|W_{Y}(t,s)|\right)$
@@
$=$ $\frac{2r}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\textup{exp}\left(-\frac{r^{2}+|W_{f}(t,s)|^{2}}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\right)$
$\times I_{0}\left(\frac{2|W_{f}(t,s)|r}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\right).$
@@
$p\left(r_{2};|W_{Y}(t_{2},s_{2})|\mid W_{\Phi}(t_{1},s_{1})\right)$
$=$ $p_{\textup{Rice}}\left(r_{2};m\left(W_{\Phi}(t_{1},s_{1})\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right)$
@@
$=$ $\delta\left(r_{1}-\left|W_{f}(t_{1},s_{1})+W_{\Phi}(t_{1},s_{1})\right|\right)$
$\times p_{\textup{Rice}}\left(r_{2};m\left(W_{\Phi}(t_{1},s_{1})\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right),$
@@
$=$ $\int_{\mathbb{C}}\delta\left(r_{1}-\left|W_{f}(t_{1},s_{1})+z\right|\right)p_{\textup{Rice}}\left(r_{2};m\left(z\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right)$
$\times\frac{1}{\pi\Gamma_{11}}\textup{exp}\left(-\frac{|z|^{2}}{\Gamma_{11}}\right)dz.$
@@
$=$ $\frac{r_{1}}{\pi\Gamma_{11}}\int_{0}^{2\pi}p_{\textup{Rice}}\left(r_{2};m\left(r_{1}e^{i\theta}-W_{f}(t_{1},s_{1})\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right)$
$\times\textup{exp}\left(-\frac{|r_{1}e^{i\theta}-W_{f}(t_{1},s_{1})|^{2}}{\Gamma_{11}}\right)d\theta.$

### v2

@@ 3.3. Covariance of AWT’s magnitudes and phases
Corollary 7 shows that the elements of $\mathbf{A}$ and $\mathbf{B}$ are pairwise independent. It also follows that the random vectors $\mathbf{A}$ and $\mathbf{B}$ are uncorrelated; that is, their $m\times n$ covariance matrix is the zero matrix. However, this does not imply that $\mathbf{A}$ and $\mathbf{B}$ are jointly independent because they are not Gaussian distributed. This finding can be regarded as an initial step toward understanding the quantifications of phase-amplitude coupling and phase-phase coupling [12, 13]. Choose $m=n$, times $t_{j}=t_{m+j}=j\Delta$ for all $j=1,\ldots,m$, where $\Delta$ is a discretization period, and scales so that $s_{1}=\ldots=s_{m}$ much larger than $s_{m+1}=\ldots=s_{2m}$. Under this configuration, $\mathbf{A}$ and $\mathbf{B}$ can be interpreted as the phase and amplitude, respectively, of bandpass-filtered signals. The uncorrelation between $\mathbf{A}$ and $\mathbf{B}$ when the input is pure noise highlights the need for caution when interpreting indices designed to quantify phase-amplitude and phase-phase coupling [12, 13], particularly when higher order moments are involved. A more thorough investigation of this issue will be pursued in future work.
3.4. Gaussian approximation for the discretized AWT of non-Gaussian noise
In practical applications, signals are usually observed in discrete time through a sampling process, and the noise is often non-Gaussian, nonstationary, and causal. The AWT is usually implemented via a Riemann-sum approximation, which we refer to as the discretized AWT. It is natural to ask whether the time-frequency representation generated by the discretized AWT is Gaussian, or asymptotically Gaussian in some sense, so that the developed random field results can be applied. Clearly, if the input random process is Gaussian, then its time-frequency representation determined by the discretized AWT is Gaussian. Therefore, it is less interesting to consider the Gaussian process discussed above and its discretization. The situation becomes interesting when the input random process is non-Gaussian. Since the AWT is defined through a kernel integration, it is intuitive to guess that the resulting time-frequency representation is approximately Gaussian, even if the noise is non-Gaussian but with sufficient moments control, and nonstationary but with sufficiently short-range dependence. In this section, we show that this intuition is correct under mild assumptions.
We now introduce a discretized version of the AWT together with a control of the truncation error arising from approximating the integral by a finite Riemann sum. Assume that a random process $\{\epsilon(t)\mid t\in\mathbb{R}\}$ is observed only at the discrete time points $\{t_{j}=j/\sqrt{n}\mid j\in\mathbb{Z}\}$, where $\sqrt{n}$ represents the sampling rate on the time axis. If the sample paths of $\{\epsilon(t)\mid t\in\mathbb{R}\}$ are continuous, the AWT of $\epsilon$ can be approximated by the following Riemann sum:
$W_{\epsilon}(t_{i},s)\approx$ $\frac{1}{\sqrt{n}}\underset{j\in\mathbb{Z}}{\sum}\epsilon(t_{j})\frac{1}{\sqrt{s}}\overline{\psi\left(\frac{t_{j}-t_{i}}{s}\right)}$
$=$ $\frac{1}{\sqrt{n}}\underset{j\in\mathbb{Z}}{\sum}\epsilon_{j}\frac{1}{\sqrt{s}}\overline{\psi\left(\frac{j-i}{\sqrt{n}s}\right)}$
(46) $=$ $\frac{1}{\sqrt{n}}\underset{j\in\mathbb{Z}}{\sum}\epsilon_{i+j}\frac{1}{\sqrt{s}}\overline{\psi\left(\frac{j}{\sqrt{n}s}\right)}$
for $i\in\mathbb{Z}$ and $s>0$, where $\epsilon_{j}=\epsilon(t_{j})$. Let the scales of interest be $0<s_{1}<s_{2}<\cdots<s_{d}$, where $d\in\mathbb{N}$. Let [A,A][-A,A] be an interval outside of which $|\psi|$ is negligibly small, where $A>0$. The right-hand side of (46) can be rewritten as follows
(47) $W_{\epsilon}(t_{i},s_{\ell})\approx$ $\frac{1}{\sqrt{n}}\underset{j\in D^{(\ell)}}{\sum}\epsilon_{i+j}\frac{1}{\sqrt{s_{\ell}}}\overline{\psi\left(\frac{j}{\sqrt{n}s_{\ell}}\right)}+R_{A},$
where
$D^{(\ell)}=\left[-\lceil\sqrt{n}As_{\ell}\rceil,\lceil\sqrt{n}As_{\ell}\rceil\right]\cap\mathbb{Z},$
$\lceil\sqrt{n}As_{\ell}\rceil=\min\left\{k\in\mathbb{Z}\mid k\geq\sqrt{n}As_{\ell}\right\}$, and
$R_{A}=\frac{1}{\sqrt{n}}\underset{j\in\mathbb{Z}\setminus D^{(\ell)}}{\sum}\epsilon_{i+j}\frac{1}{\sqrt{s_{\ell}}}\overline{\psi\left(\frac{j}{\sqrt{n}s_{\ell}}\right)}.$
Note that
$\mathbb{E}[|R_{A}|]\leq$ $\sqrt{s_{\ell}}\left(\underset{k\in\mathbb{Z}}{\max}\ \mathbb{E}\left[|\epsilon_{k}|\right]\right)\left(\frac{1}{\sqrt{n}s_{\ell}}\underset{\mathbb{Z}\setminus D^{(\ell)}}{\sum}\left|\psi\left(\frac{j}{\sqrt{n}s_{\ell}}\right)\right|\right).$
Because $|\psi|\in L^{1}$,
$\frac{1}{\sqrt{n}s_{\ell}}\underset{j\in\mathbb{Z}\setminus D^{(\ell)}}{\sum}\left|\psi\left(\frac{j}{\sqrt{n}s_{\ell}}\right)\right|\rightarrow\int_{\mathbb{R}\setminus[-A,A]}\left|\psi\left(x\right)\right|dx$
when the sampling rate $\sqrt{n}$ tends to infinity. It follows that $\mathbb{E}[|R_{A}|]$ converges to zero as $A\rightarrow\infty$ if $\max_{k\in\mathbb{Z}}\ \mathbb{E}\left[|\epsilon_{k}|\right]<\infty$. Motivated by (47) and the above estimate of the remainder term $R_{A}$, we consider a discretized AWT defined by
(48) $\widetilde{W}_{\epsilon}(t_{i},s_{\ell})=\frac{1}{n^{1/4}}\underset{j\in D^{(\ell)}}{\sum}\epsilon_{i+j}\overline{\psi_{j}^{(\ell)}},$
where
(49) $\psi_{j}^{(\ell)}=\frac{1}{n^{1/4}s_{\ell}^{1/2}}\psi\left(\frac{j}{\sqrt{n}s_{\ell}}\right).$
Assumption 3.
The sequence $\{\epsilon_{j}\}_{j\in\mathbb{Z}}$ admits the causal representation [32]
(50) $\epsilon_{j}=H_{j}(\mathcal{F}_{j}),$
where $\mathcal{F}_{j}=(\ldots,e_{j-2},e_{j-1},e_{j})$, $\{e_{j}\}_{j\in\mathbb{Z}}$ is a sequence of i.i.d. random variables, and $H_{j}:\mathbb{R}^{\infty}\mapsto\mathbb{R}$ is a measurable function describing the causal filtration mechanism.
The function $H_{j}$ models the noise generation process. When $H_{j}=H$ for a fixed function $H$, the noise generation process is fixed, and the random sequence $\{\epsilon_{j}\}_{j\in\mathbb{Z}}$ is stationary. In practice, this noise generation process is meaningless without further assumptions. We focus on two controls: moment conditions and temporal dependence. We follow the dependence measure proposed in [32] to control the dependence structure of $\{\epsilon_{j}\}_{j\in\mathbb{Z}}$. Let $\{\widehat{e}_{j}\}_{j\in\mathbb{Z}}$ be an independent copy of $\{e_{j}\}_{j\in\mathbb{Z}}$. For $k\in\mathbb{N}\cup\{0\}$, denote
$\mathcal{F}_{j,j-k}=(\ldots,e_{j-k-1},\widehat{e}_{j-k},e_{j-k+1},\ldots,e_{j-2},e_{j-1},e_{j}).$
Assumption 4.
Assume that $\sup_{j\in\mathbb{Z}}\ \mathbb{E}[|\epsilon_{j}|^{p}]<\infty$ for some $p>2$, and that
$\sup_{j\in\mathbb{Z}}\ \left(\mathbb{E}\left[\left|H_{j}(\mathcal{F}_{j})-H_{j}(\mathcal{F}_{j,j-k})\right|^{p}\right]\right)^{1/p}=\mathcal{O}\left((k+1)^{-(\chi+1)}(\log(k+1))^{-a}\right)$
for some constants $\chi>1$ and $a>\sqrt{\chi}+1$; that is, the influence of the $k$-step historical input $e_{j-k}$ on the output $\epsilon_{j}$ decays polynomially.
Define a function $G:(2,\infty)\times(1,\infty)\mapsto(0,\infty)$ by [33]
$G(p,\chi)=\frac{1}{1-\frac{1}{\kappa}-\frac{1}{p}}\left(\frac{1}{2}-\frac{1}{p}-\frac{1}{\kappa}+\frac{2}{p\kappa}\right)$
where
$\frac{1}{\kappa}=\max\Biggl\{\frac{1}{p},\;\frac{1}{\sqrt{\chi+1}\,p}$ $+\left(\frac{1}{2}-\frac{1}{\sqrt{\chi+1}\,p}\right)\max\left\{\frac{2}{\sqrt{\chi+1}\,p},\;\frac{1}{\chi}\!\left(\frac{1}{2}-\frac{1}{p}\right)\right\}\Biggr\}.$
Note that $G(p,\chi)\in(0,1/2-1/p)$ for $p>2$ and $\chi>1$. The graph of $G(p,\cdot)$ for several values of $p$ is shown in Figure 8.
Figure 8. Graph of function $G(p,\cdot)$ for several values of $p$.
[…]
$\pi\!\left(\nu_{1},\nu_{2}\right)=\inf\left\{\varepsilon>0:\;\begin{aligned} &\nu_{1}(B)\leq\nu_{2}(B^{\varepsilon})+\varepsilon,\\ &\nu_{2}(B)\leq\nu_{1}(B^{\varepsilon})+\varepsilon,\\ &\forall\,B\in\mathcal{B}(\mathbb{C}^{n\times d})\end{aligned}\right\},$
where $\mathcal{B}(\mathbb{C}^{n\times d})$ denotes the Borel $\sigma$-algebra on $\mathbb{C}^{n\times d}$, and $B^{\varepsilon}=\{x\in\mathbb{C}^{n\times d}\mid\underset{x^{\prime}\in B}{\inf}|x-x^{\prime}|_{\infty}<\varepsilon\}$ is the $\varepsilon$-enlargement of $B$ with respect to the supremum-norm metric. We have the following Gaussian approximation result for the discretized AWT.
Proposition 8.
[…]
(51) $\pi(\nu_{\epsilon}^{(n)},\nu_{z}^{(n)})=o\left(n^{-G(p,\chi)/3}(\log n)^{4/3+\delta}\right)$
as $n$ tends to infinity, where $\delta>0$ is arbitrary, and $\widetilde{W}_{z}$ is defined in the same way as in (48) but with the sequence $\{\epsilon_{j}\}_{j\in\mathbb{Z}}$ replaced by the Gaussian sequence $\{z_{j}\}_{j\in\in\mathbb{Z}}$.
The proof is provided in Section 5.15. Proposition 8 shows that, under the moment and dependence control assumptions, the distribution of the discretized AWT of a causal non-Gaussian time series, which can be nonstationary, can be approximated by that of a suitably constructed Gaussian sequence asymptotically as $n\to\infty$, with an explicit convergence rate as the sample size increases. The appearance of the factor $\sqrt{s_{\ell}}$ in front of $\widetilde{W}_{\epsilon}(t_{i},s_{\ell})$ and $\widetilde{W}_{z}(t_{i},s_{\ell})$ shows that the Gaussian approximation of the distribution of $\widetilde{W}_{\epsilon}(t_{i},s_{\ell})$ by that of $\widetilde{W}_{z}(t_{i},s_{\ell})$ becomes increasingly accurate as the scale $s_{\ell}$ grows. A scale-dependent improvement in Gaussian approximation has also been observed in [34].
To take a closer look at the bound (51), for $i=1,\ldots,n$, define the $d$-dimensional vector
$\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})=\left[\widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{1})\ \widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{2})\ \cdots\ \widetilde{W}_{\widehat{\epsilon}}(t_{i},s_{d})\right]\,$,
where $\{\widehat{\epsilon}_{j}\}_{j\in\mathbb{Z}}$ is a random sequence with the same distribution as $\{\epsilon_{j}\}_{j\in\mathbb{Z}}$. The vector $\widetilde{\mathbf{W}}_{z}(t_{i})$ is defined analogously. In parallel with [33, Theorem 3.2], if the scale values are bounded, that is, $\{s_{\ell}\}_{1\leq\ell\leq d}\subset[s_{*},s^{*}]$ for some constants $s_{*}$ and $s^{*}\in(0,\infty)$, then the Euclidean norm in $\mathbb{R}^{d}$ of $\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})$ admits the following estimate
$\mathbb{E}\left[\underset{1\leq i\leq n}{\max}\left|\widetilde{\mathbf{W}}_{\widehat{\epsilon}}(t_{i})-\widetilde{\mathbf{W}}_{z}(t_{i})\right|^{2}\right]=\mathcal{O}\left(dn^{-G(p,\chi)}\left(\log\left(n\right)\right)^{4}\right),$
which means that the approximation error converges to zero not only in the case of a fixed number $d$ of scales, but also when $d$ grows with $n$ according to $d=n^{G(p,\chi)-\vartheta}$ for some $\vartheta>0$. When $p,\chi\to\infty$ so that $G(p,\chi)\to 1/2$, the number of scales we can handle is $d=n^{1/2-\vartheta}$ for some $\vartheta>0$. Note that this rate is faster than that of discretized short-time Fourier transform shown in [33, Theorem 3.2]. See Remark 4 at the end of Section 5.15 for details.
We shall mention that Proposition 8 provides a link between the discretized AWT of non-Gaussian noise and the continuous-time AWT of Gaussian noise, in a manner analogous to the relationship between stochastic differential equations and their discretized stochastic difference equations driven by non-Gaussian innovations [35, 36]. Future work could focus on achieving a better convergence rate and extending the results above to more general noise models. Since nonstationary and non-Gaussian noise is everywhere, we also need to study what is the suitable continuous model whose discretization satisfies the causal condition and Assumption 4, and establish the relevant results.
@@ 4. Conclusion
This paper presented a statistical framework for analyzing the magnitude and phase of the AWT of noisy signals. Modeling the observation as a deterministic signal corrupted by stationary Gaussian noise, we derived joint and marginal probability density functions for the magnitude and phase in the general and noise-only settings. The analysis yielded SNR-dependent concentration bounds, explicit covariance expressions, and independence results, offering mathematical foundations for their distributional and dependence structures. We further connected these results to practical tasks by characterizing ridge detection error probabilities, the regularity of scalogram contours and the relationship between the magnitude and phase. In addition, we extended the framework to non-Gaussian noise by establishing a Gaussian approximation for the discretized AWT of weakly dependent noise sequences. The results lay the groundwork for developing rigorous AWT-based algorithms and wavelet-domain statistical inference.
@@ 5.1. Proof of Lemma 1
$(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}\Gamma(\ell,\ell^{\prime})=$ $(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}\mathbb{E}\left[W_{\Phi}(t_{\ell},s_{\ell})\overline{W_{\Phi}(t_{\ell^{\prime}},s_{\ell^{\prime}})}\right]$
@@
(54) $=$ $\int_{\mathbb{R}}\int_{\mathbb{R}}e^{it_{\ell}\lambda}\overline{\widehat{\psi}(s_{\ell}\lambda)}e^{-it_{\ell^{\prime}}\lambda^{\prime}}\widehat{\psi}(s_{\ell^{\prime}}\lambda^{\prime})\mathbb{E}\left[Z_{F}(d\lambda)\overline{Z_{F}(d\lambda^{\prime})}\right].$
@@
$(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}C_{\ell\ell^{\prime}}=$ $(s_{\ell}s_{\ell^{\prime}})^{-\frac{1}{2}}\mathbb{E}\left[W_{\Phi}(t_{\ell},s_{\ell})W_{\Phi}(t_{\ell^{\prime}},s_{\ell^{\prime}})\right]$
@@
(55) $=$ $\int_{\mathbb{R}}\int_{\mathbb{R}}e^{it_{\ell}\lambda}\overline{\widehat{\psi}(s_{\ell}\lambda)}e^{it_{\ell^{\prime}}\lambda^{\prime}}\overline{\widehat{\psi}(s_{\ell^{\prime}}\lambda^{\prime})}\mathbb{E}\left[Z_{F}(d\lambda)Z_{F}(d\lambda^{\prime})\right].$
@@ 5.2. Proof of Proposition 1
$p\left(r;|W_{Y}(t,s)|\right)=$ $\frac{r}{\pi\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\int_{-\pi}^{\pi}\textup{exp}\left(-\frac{\left|re^{i\theta}-W_{f}(t,s)\right|^{2}}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\right)d\theta$
@@
$p\left(r;|W_{Y}(t,s)|\right)=$ $\frac{r}{\pi\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\textup{exp}\left(-\frac{r^{2}+|W_{f}(t,s)|^{2}}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\right)$
@@
$=$ $\frac{2r}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\textup{exp}\left(-\frac{r^{2}+|W_{f}(t,s)|^{2}}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\right)I_{0}\left(\frac{2|W_{f}(t,s)|r}{\mathbb{E}[|W_{\Phi}(t,s)|^{2}]}\right).$
@@
$p\left(r_{2};|W_{Y}(t_{2},s_{2})|\mid W_{\Phi}(t_{1},s_{1})\right)=p_{\textup{Rice}}\left(r_{2};m\left(W_{\Phi}(t_{1},s_{1})\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right)$
@@
$=$ $\delta\left(r_{1}-\left|W_{f}(t_{1},s_{1})+W_{\Phi}(t_{1},s_{1})\right|\right)p_{\textup{Rice}}\left(r_{2};m\left(W_{\Phi}(t_{1},s_{1})\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right),$
@@
$=$ $\int_{\mathbb{C}}\delta\left(r_{1}-\left|W_{f}(t_{1},s_{1})+z\right|\right)p_{\textup{Rice}}\left(r_{2};m\left(z\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right)\frac{1}{\pi\Gamma_{11}}\textup{exp}\left(-\frac{|z|^{2}}{\Gamma_{11}}\right)dz.$
@@
$=$ $\frac{r_{1}}{\pi\Gamma_{11}}\int_{0}^{2\pi}\hskip-5.69046ptp_{\textup{Rice}}\left(r_{2};m\left(r_{1}e^{i\theta}-W_{f}(t_{1},s_{1})\right),\frac{1}{2}\mathbb{E}[|V|^{2}]\right)\textup{exp}\left(-\frac{|r_{1}e^{i\theta}-W_{f}(t_{1},s_{1})|^{2}}{\Gamma_{11}}\right)d\theta.$